An installer can carpet a room in 3 hrs, an assistant needs 4.5 hours. If the assistant helps for one hour and then is called away, how long will it take the installer to finish

Answer by stanbon(75887)

An installer can carpet a room in 3 hrs, an assistant needs 4.5 hours. If the assistant helps for one hour and then is called away, how long will it take the installer to finish
====
Installer rate = 1/3 job/hr
Assistant rate = 1/(9/2) = 2/9 job/hr
-----
Equations:
work + work = job done
1(2/9) + (1/3)x = 1
----
(1/3)x = 7/9
---
x = 7/3 = 2 1/3 hrs
=========================
Cheers,
Stan H.
===========